# Beyond Reactivity: Measuring Proactive Problem Solving in LLM Agents

### v2

@@ 8 Limitations and Future Work
Appendix A Appendix A: Baseline Implementations
@@ A.3 ReWOO
Appendix B Appendix B: Data Generation Prompts
@@
B.1 World Model Generation Prompts
@@
B.1.1 Generate Actions for Bottleneck
@@ 4. Total number of actions equals {{ num_actions }}
Listing 1: generate_actions_for_bottleneck.j2
B.1.2 Generate Organization Structure
@@ 5. Key processes and workflows
Listing 2: generate_org_structure.j2
B.1.3 Generate Personal Context
@@ 5. Tools and systems they use
Listing 3: generate_personal_context.j2
B.1.4 Generate Relationships
@@ 5. Current status of the relationship
Listing 4: generate_relationships.j2
B.2 Bottleneck Injection Prompts
@@
B.2.1 Generate Individual Bottleneck
@@ 5. Is solvable through proactive action
Listing 5: generate_bottleneck.j2
B.2.2 Generate Bottlenecks Batch
@@ 5. Is solvable through action
Listing 6: generate_bottlenecks_batch.j2
B.3 Checklist Generation Prompts
@@
B.3.1 Three-Step Checklist
Usage Context This prompt generates the core three-step checklist (Search → Identification → Task Selection) for evaluating agent performance on a specific bottleneck. It’s the primary evaluation structure.
@@ B.3.1 Three-Step Checklist
Listing 7: three_step_checklist.j2
B.4 True Positive Generation Prompts
@@
B.4.1 Plan Evidence Distribution
@@ 4. Make evidence discoverable but not too obvious
Listing 8: plan_evidence_distribution.j2
B.4.2 Generate Email Evidence
@@ 6. Sounds natural and professional
Listing 9: generate_email_evidence.j2
B.4.3 Generate Calendar Evidence
@@ 5. Completely avoids other bottlenecks
Listing 10: generate_calendar_evidence.j2
B.4.4 Generate Document Evidence
@@ 7. Never mentions other bottlenecks
Listing 11: generate_document_evidence.j2
B.4.5 Generate Dynamic Sources
@@ 4. Why its relevant to this bottleneck
Listing 12: generate_dynamic_sources.j2
B.4.6 Review Evidence
@@ 5. Search queries that would find this evidence
Listing 13: review_evidence.j2
B.5 Distractor Generation Prompts
@@
B.5.1 Generate Email Distractors
@@ 6. Length: 150-300 words each
Listing 14: generate_email_distractors.j2
B.5.2 Generate Calendar Distractors
@@ 6. Reviews or retrospectives
Listing 15: generate_calendar_distractors.j2
B.5.3 Generate Document Distractors
@@ 6. Policy documents
Listing 16: generate_document_distractors.j2
B.5.4 Generate Natural Distractor
@@ 5. Matches typical {{ kind }} format and style
Listing 17: generate_natural_distractor.j2
B.5.5 Enhance Distractor
@@ 5. Make it harder to distinguish from true positives
Listing 18: enhance_distractor.j2
B.5.6 Counterfactual Email Template
Usage Context This prompt generates counterfactual scenarios - communications about what could have happened but didn’t, useful for testing agent reasoning about hypotheticals.
Create an email discussing a hypothetical or counterfactual scenario.
CONTEXT:
- Sender: {{ sender_name }}
- Recipient: {{ recipient_name }}
- Topic: What might have happened if a different decision was made
The email should:
1. Discuss alternative scenarios or outcomes
2. Be clearly hypothetical/counterfactual
3. Not provide evidence for actual bottlenecks
4. Maintain professional tone
Example: "If we had chosen vendor B instead of vendor A..."
Listing 19: counterfactual_min_email.j2
B.5.7 Entity Swap Template
Usage Context This prompt creates distractors by swapping entities (people, projects, systems) to create plausible but incorrect variations of real situations.
Create a distractor by swapping key entities in a work scenario.
ORIGINAL ENTITIES:
- Person: {{ original_person }}
- Project: {{ original_project }}
- System: {{ original_system }}
SWAP TO:
- Person: {{ swap_person }}
- Project: {{ swap_project }}
- System: {{ swap_system }}
Generate a {{ content_type }} that uses the swapped entities in a realistic but unrelated context.
Listing 20: entity_swap.j2
B.5.8 Time Scope Template
Usage Context This prompt generates distractors that reference different time periods, useful for testing whether agents can distinguish current issues from past or future ones.
Generate a {{ content_type }} that references a different time period.
ACTUAL TIMEFRAME: {{ actual_timeframe }}
DISTRACTOR TIMEFRAME: {{ distractor_timeframe }}
Create content that:
1. Clearly references the distractor timeframe
2. Discusses similar topics but in a different temporal context
3. Could be confused with current issues if not reading carefully
4. Maintains realism for that time period
Listing 21: time_scope.j2
B.5.9 Template Clone Document
Usage Context This prompt creates documents that follow the same template or format as important documents but contain different, unrelated content.
Create a document that follows the same template as important documents but with unrelated content.
TEMPLATE TYPE: {{ template_type }}
TYPICAL SECTIONS: {{ sections | join(’, ’) }}
Generate a document that:
1. Follows the exact same structure
2. Contains completely different, unrelated content
3. Maintains professional appearance
4. Could be mistaken for an important document if not read carefully
Listing 22: template_clone_doc.j2
B.5.10 Generic Counterfactual
Usage Context This prompt generates generic counterfactual content that discusses alternatives, hypotheticals, or what-if scenarios unrelated to actual bottlenecks.
Generate {{ kind }} content discussing counterfactual scenarios for {{ world_model.persona_full_name }}.
CONTEXT:
- Role: {{ world_model.persona_occupation }}
- Department: {{ world_model.organizational_structure.department }}
Create content about:
- Alternative approaches that werent taken
- Hypothetical scenarios
- Lessons learned from paths not chosen
- What-if analyses
Ensure the counterfactual discussion:
1. Is clearly marked as hypothetical
2. Doesnt relate to actual bottlenecks
3. Provides interesting but irrelevant information
4. Maintains professional tone
Listing 23: generic_counterfactual.j2
B.6 LLM-as-Judge Evaluation Prompts

### v3

@@ 6 Limitations and Future Work
Appendix A Appendix A: Related Work
Emerging proactive agents face the challenge of anticipating latent user goals from partial observations and executing actions without explicit instruction. Recent work explores intent inference from behavioral patterns (Zhao et al., 2025), continuous insight generation from data streams (Yang et al., 2025a), and context-aware action generation (Shaikh et al., 2025). However, these systems lack systematic evaluation frameworks that assess end-to-end proactive capabilities.
Agent Benchmarking: Existing benchmarks predominantly evaluate reactive systems: SWE-bench (Jimenez et al., 2024) for software engineering; ToolBench (Qin et al., 2023) for API calling; and GAIA (Mialon et al., 2023) for multi-hop reasoning. Recent proactivity benchmarks like ProactiveVideoQA (Wang et al., 2025) and ProCIS (Samarinas and Zamani, 2024) begin addressing this gap but remain limited to conversational actions.
The Proactive Agent framework Lu et al. (2024) serves as our nearest counterpart, contributing a benchmark for annotated proactive suggestions in coding, writing, and everyday tasks. We build upon this work by benchmarking proactivity in long-horizon and multi-document datastores, incorporating search to uncover and address opportunities for assistance.
The above-mentioned works do not decompose proactivity into constituent capabilities or test comprehensive task execution across a multitude of sources and extended time horizons, motivating the systematic approach found in PROBE .
Appendix B Appendix B: Baseline Implementations
@@ B.3 ReWOO
Appendix C Appendix C: Example World Model and Bottleneck
This Appendix provides a complete example of a world model and a bottleneck used in our benchmark. The example demonstrates the persona context, relationships, and available actions that the agent must reason over.
C.1 World Model
World Model Example The following JSON structure represents a complete world model instance from our benchmark dataset, illustrating the personalized level of detail and complexity agents must navigate in PROBE.
"world_model": {
"persona_id": "connor_smith",
"persona_full_name": "Connor Smith",
"persona_occupation": "Financial Services Lead Advisory, PwC UK",
"persona_about": "A Business graduate with a strong interest in Finance currently working for PwC within the Corporate Finance practice. ACA exam qualified with ICAEW, all first time passes. A Manager in the Financial Services Lead Advisory team within PwC UK. The FSLA team combines deep industry knowledge and transaction experience to advise on all aspects of corporate finance M&A and, through the Portfolio Advisory Group, credit related transactions. The team leverages more than 6,000 PwC people in the U.K. and 60,000 PwC people globally who work day in day out with financial services clients.",
"relationships": [
"name": "Emily Patel",
"type": "manager",
"department": "Financial Services Lead Advisory",
"relationship_context": "Direct line manager overseeing Connors progression from Senior Associate to Manager, providing strategic direction on major M&A deals and regular feedback on client engagement style."
"name": "James Thornton",
"type": "colleague",
"department": "Portfolio Advisory Group",
"relationship_context": "Collaborate intensively on distressed credit portfolio transactions, often negotiating differing approaches to valuation methodologies but ultimately delivering cohesive client recommendations."
"name": "Priya Malhotra",
"type": "direct_report",
"department": "Financial Services Lead Advisory",
"relationship_context": "Recently promoted Senior Associate who Connor is mentoring through her first lead on a mid-market bank acquisition, balancing support with constructive critique."
"name": "Olivia Green",
"type": "client",
"department": null,
"relationship_context": "Senior Corporate Development Manager at a challenger bank, relies on Connors team for buy-side advisory; the relationship is trusted but currently tense due to a delayed regulatory approval."
"name": "Michael Zhang",
"type": "stakeholder",
"department": "Risk Assurance",
"relationship_context": "Internal stakeholder who reviews risk elements of Connors transactions; their differing perspectives on risk appetite have led to constructive debates and enhanced deal structuring."
"name": "Sophie Laurent",
"type": "vendor",
"department": null,
"relationship_context": "Relationship manager at a third-party data analytics firm providing due diligence support, with whom Connor negotiates service level agreements and escalates data quality issues."
"name": "Rebecca Yates",
"type": "stakeholder",
"department": "Compliance",
"relationship_context": "Works closely with Connor to ensure all deal processes meet FCA regulations, occasionally challenging his teams interpretation of compliance requirements."
"name": "Tomás Ortega",
"type": "client",
"department": null,
"relationship_context": "Head of M&A at a European insurance group, Connor manages international communications and project delivery, navigating cultural nuances and tight deadlines."
"name": "Grace Lin",
"type": "mentee",
"department": "Financial Services Lead Advisory",
"relationship_context": "Graduate Analyst whom Connor is coaching on financial modeling; their regular sessions have fostered Graces rapid development, though she sometimes struggles with client-facing confidence."
"name": "Andrew Williams",
"type": "cross_functional_partner",
"department": "Technology Consulting",
"relationship_context": "Collaborate on digital transformation due diligence for fintech targets, balancing differing priorities between operational efficiency and deal speed."
"name": "Marta Kowalska",
"type": "vendor",
"department": null,
"relationship_context": "Director at a legal advisory firm providing transaction legal due diligence; Connor manages the engagement and ensures alignment on documentation timetables."
"name": "Ben Carter",
"type": "direct_report",
"department": "Financial Services Lead Advisory",
"relationship_context": "Associate who handles initial deal screening and supports Connor on market research, recently needing extra guidance due to a challenging workload."
"name": "Hannah Robinson",
"type": "colleague",
"department": "Valuations",
"relationship_context": "Works with Connor on valuations for complex asset portfolios; their differing technical approaches often lead to productive, data-driven debates."
"name": "Samir Ahmed",
"type": "stakeholder",
"department": "Human Capital",
"relationship_context": "Partners with Connors team for talent allocation on high-value deals; occasionally tensions arise during peak periods over resource prioritization."
"name": "Julia Becker",
"type": "client",
"department": null,
"relationship_context": "Private equity partner who values Connors candid market assessments, but has recently challenged his teams fee structure on a major divestment mandate."
"name": "Freddie White",
"type": "mentee",
"department": "Financial Services Lead Advisory",
"relationship_context": "Early-career team member Connor supports through the ICAEW qualification process, providing both technical guidance and emotional support during exam periods."
"name": "Isabelle Dubois",
"type": "cross_functional_partner",
"department": "Sustainability & ESG Advisory",
"relationship_context": "Works with Connor to integrate ESG considerations into financial advisory pitches, sometimes pushing Connor to adapt more progressive frameworks for skeptical clients."
"personal_context": {
"work_style": "Connor operates in a high-pressure, fast-paced environment, balancing multiple M&A transaction pipelines with ongoing client advisory work. He prioritizes structured project management, but frequently adapts to shifting client demands and tight regulatory timelines. He leverages both collaborative teamwork and independent analysis, often working across time zones and balancing strategic oversight with detailed financial diligence.",
"communication_preferences": [
"Email for formal updates and transaction documentation",
"Video calls for deal negotiation and client relationship management",
"Instant messaging (e.g., Microsoft Teams) for quick internal queries and urgent project coordination",
"Phone calls for time-sensitive decisions or when clarity is critical",
"Face-to-face meetings for complex client presentations and team alignment sessions"
"current_priorities": [
"Managing multiple concurrent M&A transactions with overlapping deadlines",
"Coordinating due diligence efforts across PwC teams and external stakeholders",
"Ensuring regulatory compliance and risk mitigation within all live deals",
"Developing junior team members while maintaining project delivery standards",
"Enhancing client relationships and identifying cross-sell opportunities",
"Integrating Portfolio Advisory Group insights into standard deal processes",
"Meeting internal performance and billing targets while navigating resource constraints"
"pain_points": [
"Juggling conflicting deadlines across several high-stakes projects",
"Difficulty in accessing timely inputs from cross-functional or global teams",
"Managing client expectations amid rapidly changing financial markets",
"Keeping up with evolving regulatory requirements and ensuring compliance",
"Limited time for proactive client development due to reactive workload",
"Onboarding and mentoring new staff without compromising on project quality"
"available_actions": [
"id": "conduct_peer_review_of_draft_report",
"type": "conduct_review",
"description": "Carry out a peer review of a draft report to ensure quality and compliance standards.",
"constraints": [
"peer_reviewer_assigned",
"review_criteria_defined"
"params_schema": {
"required": [
"document_id",
"reviewer",
"review_feedback"
"id": "create_briefing_note_on_sector_updates",
"type": "create_document",
"description": "Develop a briefing note summarizing key sector updates for internal awareness.",
"constraints": [
"sector_data_compiled",
"briefing_template_used"
"params_schema": {
"required": [
"briefing_date",
"sector",
"summary_points"
"id": "review_and_approve_time_entries",
"type": "update_task_status",
"description": "Update the approval status of submitted time entries for the previous month.",
"constraints": [
"time_entries_submitted",
"review_period_open"
"params_schema": {
"required": [
"timesheet_id",
"approval_status",
"review_comments"
"id": "finalize_presentation_for_external_workshop",
"type": "update_document",
"description": "Make final adjustments to the presentation slides for an upcoming external workshop.",
"constraints": [
"slide_content_approved",
"branding_guidelines_followed"
"params_schema": {
"required": [
"document_id",
"sections_to_edit",
"visual_elements"
"id": "initiate_internal_knowledge_share",
"type": "schedule_meeting",
"description": "Arrange a knowledge-sharing session for the team to exchange best practices on due diligence processes.",
"constraints": [
"subject_matter_experts_available",
"content_outline_prepared"
"params_schema": {
"required": [
"attendees",
"date",
"time",
"session_topic"
"id": "send_reminder_for_time_allocation_reconciliation",
"type": "send_email",
"description": "Send a reminder to the operations team to expedite reconciliation of time allocation data in the finance system.",
"constraints": [
"prior_request_sent",
"pending_reconciliation"
"params_schema": {
"required": [
"recipients",
"subject",
"reminder_message"
"id": "initiate_annual_training_enrollment",
"type": "create_task",
"description": "Create a task to enroll the team in mandatory annual training sessions.",
"constraints": [
"training_schedule_available",
"team_member_list_current"
"params_schema": {
"required": [
"training_type",
"enrollment_deadline",
"participant_list"
"id": "organize_quarterly_team_checkin",
"type": "schedule_meeting",
"description": "Schedule a quarterly meeting for the team to discuss ongoing initiatives and operational improvements.",
"constraints": [
"team_calendar_availability",
"meeting_room_availability"
"params_schema": {
"required": [
"attendees",
"date",
"time",
"agenda"
"id": "submit_request_for_client_feedback",
"type": "request_approval",
"description": "Submit a formal request for client feedback on recently delivered project outputs.",
"constraints": [
"project_deliverables_completed",
"client_point_of_contact_confirmed"
"params_schema": {
"required": [
"client_id",
"project_id",
"feedback_request_message"
"id": "initiate_information_request_to_central_finance_team",
"type": "send_email",
"description": "Send a follow-up email to the central finance group to request updated reference data required for current project models.",
"constraints": [
"previous_request_no_response",
"project_deadlines_approaching"
"params_schema": {
"required": [
"recipients",
"subject",
"message_body"
"id": "reply_to_team_query_on_policy_update",
"type": "reply_email",
"description": "Reply to a team members email regarding recent policy updates and clarify outstanding questions.",
"constraints": [
"policy_documentation_on_hand",
"clarification_needed"
"params_schema": {
"required": [
"thread_id",
"response_content"
"id": "cancel_redundant_calendar_event",
"type": "cancel_meeting",
"description": "Cancel an outdated or redundant calendar event to reduce scheduling conflicts.",
"constraints": [
"event_no_longer_needed",
"invitees_notified"
"params_schema": {
"required": [
"meeting_id",
"cancellation_reason"
"id": "reschedule_internal_alignment_meeting",
"type": "reschedule_meeting",
"description": "Move the internal alignment meeting to accommodate conflicting priorities.",
"constraints": [
"attendee_availability_checked",
"calendar_invite_updated"
"params_schema": {
"required": [
"meeting_id",
"new_date",
"new_time"
"id": "prepare_meeting_agenda_for_upcoming_review",
"type": "create_meeting_agenda",
"description": "Draft an agenda for an upcoming review meeting with stakeholders.",
"constraints": [
"meeting_objectives_defined",
"stakeholder_input_collected"
"params_schema": {
"required": [
"meeting_date",
"agenda_items",
"expected_outcomes"
"id": "assign_documentation_task_to_team_member",
"type": "delegate_task",
"description": "Delegate the task of documenting process flows to a designated team member.",
"constraints": [
"team_member_availability",
"skill_relevance"
"params_schema": {
"required": [
"assignee",
"task_description",
"due_date"
"id": "request_additional_resources_for_project",
"type": "create_task",
"description": "Open a task to formally request extra resources for a high-priority project due to increased workload.",
"constraints": [
"resource_justification_prepared",
"project_timeline_affected"
"params_schema": {
"required": [
"project_id",
"resource_type",
"rationale"
"id": "compile_weekly_market_summary_report",
"type": "create_document",
"description": "Prepare a weekly summary document highlighting recent market trends for internal distribution.",
"constraints": [
"market_data_access",
"standard_formatting"
"params_schema": {
"required": [
"report_date",
"market_segments",
"key_findings"
"id": "follow_up_on_pending_client_information",
"type": "send_slack_message",
"description": "Send a message to the client liaison to check on the status of outstanding information requests.",
"constraints": [
"previous_request_sent",
"response_delay"
"params_schema": {
"required": [
"recipient",
"message_content"
"id": "send_status_update_on_deliverables",
"type": "send_email",
"description": "Send an email to project stakeholders with a status update on key deliverables and next steps.",
"constraints": [
"stakeholder_list_current",
"deliverable_status_confirmed"
"params_schema": {
"required": [
"recipients",
"subject",
"update_content"
"id": "make_call_to_confirm_external_deadlines",
"type": "make_phone_call",
"description": "Place a call to confirm upcoming external deadlines for a strategic initiative.",
"constraints": [
"contact_information_verified",
"relevant_documents_prepared"
"params_schema": {
"required": [
"contact_number",
"call_objective",
"call_notes"
"id": "escalate_technical_support_ticket",
"type": "escalate_to_manager",
"description": "Escalate an unresolved technical support ticket to the appropriate escalation channel for prioritized handling.",
"constraints": [
"ticket_open_exceeds_threshold",
"business_impact_documented"
"params_schema": {
"required": [
"ticket_id",
"escalation_reason",
"desired_resolution_timeframe"
"id": "update_internal_guidance_document",
"type": "update_document",
"description": "Revise internal guidance materials to reflect recent procedural changes.",
"constraints": [
"approval_for_edits",
"current_policy_reviewed"
"params_schema": {
"required": [
"document_id",
"sections_to_update",
"revision_notes"
"id": "share_internal_best_practices_document",
"type": "share_document",
"description": "Distribute a document outlining internal best practices to the wider department.",
"constraints": [
"document_reviewed",
"distribution_list_confirmed"
"params_schema": {
"required": [
"document_id",
"recipients",
"share_message"
"id": "update_project_milestone_tracker",
"type": "update_project_plan",
"description": "Revise the project milestone tracker to reflect recent changes in timelines.",
"constraints": [
"latest_project_updates_collected",
"tracker_accessible"
"params_schema": {
"required": [
"project_id",
"milestone_updates",
"responsible_owners"
"id": "request_access_to_new_analytics_tools",
"type": "request_access",
"description": "Submit an access request for recently deployed analytics platforms to support project work.",
"constraints": [
"manager_approval_obtained",
"tool_list_reviewed"
"params_schema": {
"required": [
"tool_name",
"user_id",
"business_justification"
"id": "draft_feedback_on_proposed_project_approach",
"type": "provide_feedback",
"description": "Provide written feedback on the proposed approach for a key project based on available guidance.",
"constraints": [
"guidance_document_accessible",
"recent_project_overview_reviewed"
"params_schema": {
"required": [
"feedback_recipient",
"project_id",
"feedback_content"
"organizational_structure": {
"company_name": "PwC UK",
"department": "Financial Services Lead Advisory (FSLA)",
"team_size": 16,
"reporting_to": "Director, Financial Services Lead Advisory",
"key_meetings": [
"Weekly FSLA team standup",
"Monday cross-sector M&A pipeline review",
"Bi-weekly Portfolio Advisory Group sync",
"Monthly Financial Services Risk Committee",
"Quarterly Executive Steering Committee",
"Client transaction strategy sessions",
"Regulatory and Compliance update",
"Global FSLA practice call",
"Cross-regional deal origination forum",
"Ad hoc project war room"
"context_difficulty": "hard"
Listing 1: Complete world model example for Connor Smith
Bottleneck Examples The following JSON structure represents a complete bottleneck from our benchmark dataset. Bottlenecks are often multifasceted and require information stemming from multiple sources.
"bottleneck": {
"description": "A critical data extract from Sophie Laurents team at DataInsight Analytics, required by the risk-adjusted NPL valuation report (needed during Mondays IC review), was delivered in an incompatible file format (SAS instead of Excel), and your IT support ticket (#PWCUK-90812) to convert the file has been open 3 days without response.",
"evidence_required": [
"document_fb96fa16",
"document_46b3383d",
"document_236c752a",
"document_6ea213cf",
"document_982d62dc"
Listing 2: Complete bottleneck example for Connor Smith
Appendix D Appendix D: Data Generation Prompts
@@
D.1 World Model Generation Prompts
@@
D.1.1 Generate Actions for Bottleneck
@@ 4. Total number of actions equals {{ num_actions }}
Listing 3: generate_actions_for_bottleneck.j2
D.1.2 Generate Organization Structure
@@ 5. Key processes and workflows
Listing 4: generate_org_structure.j2
D.1.3 Generate Personal Context
@@ 5. Tools and systems they use
Listing 5: generate_personal_context.j2
D.1.4 Generate Relationships
@@ 5. Current status of the relationship
Listing 6: generate_relationships.j2
D.2 Bottleneck Injection Prompts
@@
D.2.1 Generate Individual Bottleneck
@@ 5. Is solvable through proactive action
Listing 7: generate_bottleneck.j2
D.2.2 Generate Bottlenecks Batch
@@ 5. Is solvable through action
Listing 8: generate_bottlenecks_batch.j2
D.3 Checklist Generation Prompts
@@
D.3.1 Three-Step Checklist
Usage Context This prompt generates the core three-step checklist (Search → Identification → Task Selection) for evaluating agent performance on a specific bottleneck. We note that these steps need not happen concurrently: we only perform evaluation on the final output.
@@ D.3.1 Three-Step Checklist
Listing 9: three_step_checklist.j2
D.4 True Positive Generation Prompts
@@
D.4.1 Plan Evidence Distribution
@@ 4. Make evidence discoverable but not too obvious
Listing 10: plan_evidence_distribution.j2
D.4.2 Generate Email Evidence
@@ 6. Sounds natural and professional
Listing 11: generate_email_evidence.j2
D.4.3 Generate Calendar Evidence
@@ 5. Completely avoids other bottlenecks
Listing 12: generate_calendar_evidence.j2
D.4.4 Generate Document Evidence
@@ 7. Never mentions other bottlenecks
Listing 13: generate_document_evidence.j2
D.4.5 Generate Dynamic Sources
@@ 4. Why its relevant to this bottleneck
Listing 14: generate_dynamic_sources.j2
D.4.6 Review Evidence
@@ 5. Search queries that would find this evidence
Listing 15: review_evidence.j2
D.5 Distractor Generation Prompts
@@
D.5.1 Generate Email Distractors
@@ 6. Length: 150-300 words each
Listing 16: generate_email_distractors.j2
D.5.2 Generate Calendar Distractors
@@ 6. Reviews or retrospectives
Listing 17: generate_calendar_distractors.j2
D.5.3 Generate Document Distractors
@@ 6. Policy documents
Listing 18: generate_document_distractors.j2
D.5.4 Generate Natural Distractor
@@ 5. Matches typical {{ kind }} format and style
Listing 19: generate_natural_distractor.j2
D.5.5 Enhance Distractor
@@ 5. Make it harder to distinguish from true positives
Listing 20: enhance_distractor.j2
D.6 LLM-as-Judge Evaluation Prompts